Can you please help me with this question. I would really appreciate it. The circumference of a cylinder is 86 pi meter, and the height is 63 meters. Find the surface area and the volume of the cylinder.
i think you use the formula v=pir^2H

Answer by macston(5194)

.
PART 1: Surface Area
Circumference=(diameter)(pi) so diameter=86 meters, radius=43 meters
Surface area=(area of base x 2)+ area of side
area of base x 2=2(pi)r^2=(pi)(43m)^2=2(pi)(1849m^2)=(pi)(3698m^2)
area of side=(height)(circumference)=(63m)(86m)(pi)=(pi)(5418m^2)
Surface area=(pi)(3698m^2)+(pi)(5418m^2)=(pi)(9116m^2)
ANSWER 1: The surface area is 9116(pi)m^2.
.
PART 2: Volume
V=volume; r=radius; h=height
V=(pi)(r^2)h
V=(pi)(1849m^2)(63m)=(pi)(116487m^3)
ANSWER 2: The volume is 116487(pi)m^3